# The BTSbot-nearby discovery of SN 2024jlf: rapid, autonomous follow-up probes interaction in an 18.5 Mpc Type IIP supernova

### v1

@@ 1 Introduction
Flash features have historically been difficult to study, in large part due to their ephemeral nature. Niemela et al. (1985) published the first SN spectra exhibiting flash features, and Yaron et al. (2017) published a remarkable dataset on SN 2013fs, now a prototypical example of flashing Type II SNe (SNe II). The search for these features (“flash spectroscopy”) has produced samples of SNe II with spectra shortly after first light, including dozens of events exhibiting flash features (Khazov et al., 2016; Bruch et al., 2021, 2023; Jacobson-Galán et al., 2024b). Flash features have also been seen in SNe from H-stripped and He-stripped stars, but we discuss only SNe II here.
Bruch et al. (2023) (B23 henceforth; and preceded by Bruch et al. 2021) search for narrow He II emission in SNe II, which they use to define a SN as flashing. They find such emission in at least 30% of events in their sample and likely in more than half. They also find no significant difference in the light curve properties (rise time, peak magnitude, color at peak) of flashing and non-flashing events. They quantify the duration of flash features () as the duration for which narrow He II emission is visible. The events in their sample typically have days, and they find tentative evidence for an additional population of rare events with long-lived flash features ( days).
@@
One such tool is BTSbot (Rehemtulla et al., 2024a), a machine learning (ML) model for autonomously identifying bright () extragalactic transients in the ZTF data stream. BTSbot was designed to select targets for the ZTF Bright Transient Survey (BTS; Fremling et al., 2020; Perley et al., 2020; Rehemtulla et al., 2024a), which endeavors to spectroscopically classify all such sources. BTSbot was deployed into ZTF and BTS operations in 2023 October with the ability to autonomously trigger follow-up to the SED Machine (SEDM; Blagorodnova et al., 2018; Kim et al., 2022). Since then, it has sent 1000 automated spectroscopic follow-up requests. Close monitoring of triggers in the first eight months of operations revealed that 96% of BTSbot identifications were of genuine extragalactic transients.
@@ 2 BTSbot-nearby
Some minor adjustments to criteria in these filters, e.g., the real/bogus score threshold using braai (Duev et al., 2019) is increased from 0.3 (in the BTS alert fitler) to 0.7. This adjustment is made to more consistently reject bogus alerts atop very bright galactic nuclei which are often misclassified as real by braai. The moving source rejection criterion now also considers alert packet forced photometry provided by the Infrared Processing and Analysis Center (IPAC; Masci et al. 2023), which allows certain sources to satisfy the criterion sooner.
@@
The primary addition is filtering following host galaxy association. Alerts only pass the BTSbot-nearby filter if they are coincident with a NASA Extragalactic Database333https://ned.ipac.caltech.edu/ (NED) galaxy that has distance Mpc. An alert is considered coincident with a galaxy if its projected physical offset is 15 kpc, capped at a maximum angular offset of . The angular offset limit is imposed because the size of the selection region for very nearby galaxies ( Mpc) is otherwise extremely large and prone to selecting sources unrelated to the nearby galaxy. The 60 Mpc limit is motivated by the ZTF sensitivity: a 20.5 mag source at 60 Mpc corresponds to having an absolute magnitude of roughly mag, sufficient to capture SNe very early. An offset limit of 15 kpc selects the vast majority of CCSNe and SNe Ia (Fremling et al., 2020; Schulze et al., 2021) while limiting contamination from sources projected nearby but unrelated to a galaxy.
@@
Once a source is associated to a galaxy, the galaxy’s distance can be used to estimate the source’s absolute magnitude. BTSbot-nearby requires that the source have absolute magnitude in any ZTF filter of mag to reject variable stars and classical novae which are visible to ZTF in the most nearby galaxies (e.g., M31 and M33; Capaccioli et al., 1989). The source must also have at least one alert with BTSbot score greater than 0.7. This criterion is crucial for rejecting remaining contamination like non-SN stellar activity in very nearby galaxies, active galactic nuclei and cataclysmic variables projected over nearby galaxies, and more. The target must also not be coincident with a source on the Transient Name Server (TNS)444https://www.wis-tns.org that was reported more than 36 hours before the time of filtering. The final criteria considers the status of the source on ZTF’s first-party marshal Fritz555https://github.com/fritz-marshal/fritz (a SkyPortal instance; van der Walt et al., 2019; Coughlin et al., 2023); a trigger is prevented if there is (i) an existing spectroscopic follow-up request; (ii) a classification assigned; or (iii) a spectrum already present. The criteria involving the status of the source on Fritz are described in detail in Appendix B of Rehemtulla et al. (2024a). Once all criteria are met, follow-up requests can be triggered with configurable parameters.
@@ 2.1 Quality of BTSbot-nearby filtering relative to human scanning
We characterize the quality of the BTSbot-nearby filtering by computing the completeness (or “recall”) and purity (or “precision”) of triggers and comparing the latency of automated follow-up with that of traditional, human-triggered follow-up. We pass all ZTF public and partnership alerts from 1 October 2023 to 1 October 2024666The dates are determined by when the latest version of BTSbot was deployed into production and when this analysis was conducted. through the BTSbot-nearby filters and log which sources satisfy them and whether they are true positives (TPs) or FPs. False negatives (FNs) are identified by cross-matching with sources cataloged in BTS or CLU during the same time frame.
Completeness and purity are calculated identically to how they were by Rehemtulla et al. (2024a). 59 sources satisfy the BTSbot-nearby filtering in this time frame; 55 are genuine nearby transients, identified via cross-match with BTS/CLU catalogs and visual inspection: 93.2% purity. The FPs include one active galactic nucleus (AGN), one cataclysmic variable star (CV), and two distant transients that were erroneously cross-matched to a nearby galaxy. These indicate that the BTSbot filtering and the host galaxy association can still be improved. We also find that 3 genuine nearby transients which occurred during this time frame were not recovered yielding a 94.8% completeness. All of the FNs are hosted by galaxies not in NED or in a NED galaxy without either a spectroscopic redshift or a redshift-independent distance measurement.
@@
Figure 1 shows the distribution of spectroscopic follow-up latencies for nearby SNe, highlighting those involving BTSbot and that which was demonstrated by BTSbot-nearby for SN 2024jlf. Latencies for SNe with human-triggered follow-up clearly cluster around day. This matches expectations for this time-period because visual inspection and triggering of high-priority follow-up typically only occur the morning after observations. When this is the case and the follow-up instrument is co-located with the survey (e.g. ZTF and SEDM), follow-up tends to occur 24 hours after discovery. This is not always the case, however. Scanners in Europe are able to easily scan in near real-time because the Palomar night coincides with morning and afternoon in Europe. This has resulted in countless instances of early follow-up triggered by European scanners yielding valuable results (e.g., Yang et al., 2021; Sollerman et al., 2021). In fact, much of the follow-up effort supporting Bruch et al. (2021) and B23 was enabled by scanners in Europe. Because the sample used here is bounded by 1 January 2021, it does not contain any SNe in B23, but the latency statistics are representative for human scanning between January 2021 and October 2024.
@@ 3.1 Discovery and Classification
Before either TNS report, BTSbot identified ZTF24aaozxhx as a genuine bright ( mag) transient, saved it to an internal ZTF transient catalog, and triggered a high-priority follow-up request for photometry () and spectroscopy to SEDM.777In the future, BTSbot-nearby discoveries will automatically be reported to TNS to facilitate community follow-up of these important targets. The SEDM trigger was sent at 07:53:40 and SEDM began observing ZTF24aaozxhx just 7 minutes later at 08:00:51; this observation occurred only 0.7 days after first light (see Sec. 3.3.1 for details). Logs of Fritz activity, where ZTF partners scan new transient candidates, suggest that no astronomer in ZTF had viewed ZTF24aaozxhx before the spectroscopic observation had concluded. Following this discovery, we collected multi-wavelength, high-cadence follow-up observations with numerous facilities to probe the evolution of SN 2024jlf.
@@ 3.3 Photometric observations
SN 2024jlf was observed in the UV (, , ) and near-UV () with the Ultraviolet/Optical Telescope (UVOT; Roming et al., 2005) onboard the Neil Gerhels Swift Obervatory (Swift; Gehrels et al., 2004). We processed all UVOT photometry of SN 2024jlf using a Python-based wrapper888https://github.com/charliekilpatrick/Swift_host_subtraction for photometry tools in heasoft v6.28 (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc). We optimally stacked each epoch and performed aperture photometry in each band using uvotsource.
@@ 3.4 Spectroscopic observations
We conducted a thorough spectroscopic follow-up campaign to monitor the evolution of SN 2024jlf out to 94.6 days after . Spectroscopic follow-up was conducted with numerous facilities: SEDM (Blagorodnova et al., 2018; Kim et al., 2022) on the P60 telescope (Cenko et al., 2006); the Spectrograph for the Rapid Acquisition of Transients (SPRAT; Piascik et al., 2014) on the robotic 2 m Liverpool Telescope (LT; Steele et al., 2004); the Alhambra Faint Object Spectrograph and Camera (ALFOSC)999https://www.not.iac.es/instruments/alfosc/ on the 2.56 m Nordic Optical Telescope (NOT); Binospec (Fabricant et al., 2019) on the 6.5 m MMT telescope; the Low Resolution Imaging Spectrometer (LRIS; Oke et al., 1995) on the W. M. Keck Observatory’s Keck I 10 m telescope; the Keck Cosmic Web Imager (KCWI; Martin et al., 2010; Morrissey et al., 2018) on the Keck II 10 m telescope; and the Wide-Field Spectrograph (WiFeS; Dopita et al., 2007, 2010) on the Australian National University 2.3 m (ANU) Advanced Technology Telescope (ATT) at Siding Spring Observatory. We also include the FLOYDS-N spectrum uploaded to TNS by Hosseinzadeh et al. (2024).
@@
The SEDM observations are reduced by the custom pySEDM package (Rigault et al., 2019). Spectra from other facilities are reduced with standard procedures (see Matheson et al. 2000). Reduction of MMT/Binospec and Keck/LRIS spectra uses pypeit (Prochaska et al., 2020); LT/SPRAT reduction uses the custom FRODOSpec pipeline (Barnsley et al., 2012); NOT/ALFOSC reduction uses a custom pypeit-based script;101010https://gitlab.com/steveschulze/pypeit_alfosc_env Keck/KCWI reduction uses the official data reduction pipeline;111111https://kcwi-drp.readthedocs.io/ and ANU/WiFeS reduction uses PyWiFeS (see Childress et al., 2014; Carr et al., 2024).121212https://github.com/PyWiFeS/pipeline Our spectroscopic observations are summarized in Table 3.
@@
Figure 6 shows our full spectral series of SN 2024jlf. We find that SN 2024jlf exhibits features and evolution typical of SNe IIP: a mostly featureless blue continuum followed by prominent P-Cyngi lines of the hydrogen Balmer series and later of He. A broad H profile first emerges in the day SPRAT spectrum. H and bluer Balmer series lines become visible shortly afterwards; the Binospec spectrum at days shows P-Cygni profiles in H, H, H, H, and possibly H. He I is also visible in a P-Cygni profile starting in the day spectrum. Fe II absorption and the Ca near-infrared triplet are visible starting around three weeks after first light. Model spectra, shown as gray lines, are discussed in Section 4.1.
@@
Figure 7 shows the regions around features in the early spectra of SN 2024jlf which exhibit features characteristic of SNe II undergoing CSM interaction: He II 4686 and the ledge, C IV 5801, 5812, and H. We also possibly detect narrow H emission in the d FLOYDS spectrum. He II is very common in flashing SNe and its presence is often used to define a SN as flashing (e.g., Khazov et al. 2016; Bruch et al. 2021; B23). The ledge-shaped feature around 4450–4700 Å, visible in the first six epochs of spectroscopy, is typically interpreted as a sign of CSM interaction and has been seen in many interacting SNe including SN 2013fs (Yaron et al., 2017) and others in the samples of B23 and JG24. Its origin is not precisely known although it has been attributed to (i) a blend of high-ionization lines including those of carbon, nitrogen, and oxygen (e.g., Soumagnac et al., 2020); (ii) high-velocity H (e.g., Pastorello et al., 2006); and (iii) blueshifted He II 4686 (e.g., Dessart et al., 2017; Bullivant et al., 2018; Chugai & Utrobin, 2023). See Pearson et al. (2023) for further discussion and references on the ledge feature.

### v2

@@ 1 Introduction
Flash features have historically been difficult to study, in large part due to their ephemeral nature. Niemela et al. (1985) published the first SN spectra exhibiting flash features, and Yaron et al. (2017) published a remarkable dataset on SN 2013fs, now a prototypical example of flashing Type II SNe (SNe II). The search for these features (“flash spectroscopy”) has produced samples of SNe II with spectra shortly after first light, including dozens of events exhibiting flash features (Khazov et al., 2016; Bruch et al., 2021, 2023; Jacobson-Galán et al., 2024b). Short-lived narrow emission features have also been seen in SNe from stripped stars (e.g., Gangopadhyay et al., 2020, 2022; Perley et al., 2022; Davis et al., 2023; Schulze et al., 2024), but we discuss only SNe II here.
Bruch et al. (2023) (B23 henceforth; preceded by Bruch et al. 2021 and accompanied by Irani et al. 2024) search for narrow He II emission in SNe II, which they use to define a SN as flashing. They unambiguously identify such emission in at least 30% of events in their sample while more than half have moderately confident detections. They also find no significant difference in the optical light curve properties (rise time, peak magnitude, color at peak) of flashing and non-flashing events while Irani et al. (2024) find that flashers have larger initial blackbody radii and luminosities. Crucially, Irani et al. (2024) find that many SNe II do not fit shock cooling models well in the early time but do later on in their evolution, underscoring the importance of well-sampled early optical and UV light curves. B23 quantify the duration of flash features () as the duration for which narrow He II emission is visible. The events in their sample typically have days, and they find tentative evidence for an additional population of rare events with long-lived flash features ( days).
@@
One such tool is BTSbot (Rehemtulla et al., 2024a), a machine learning (ML) model for autonomously identifying bright () extragalactic transients in the ZTF data stream. BTSbot was designed to select targets for the ZTF Bright Transient Survey (BTS; Fremling et al., 2020; Perley et al., 2020; Rehemtulla et al., 2024a), which endeavors to spectroscopically classify all such sources. BTSbot was deployed into ZTF and BTS operations in 2023 October with the ability to autonomously trigger follow-up to the SED Machine (SEDM; Blagorodnova et al., 2018; Kim et al., 2022). Since then, it has sent 1200 automated spectroscopic follow-up requests, many of which are on transients it identified before human scanners in BTS. Close monitoring of triggers in the first eight months of operations revealed that 96% of BTSbot identifications were of genuine extragalactic transients, matching expected performance (Rehemtulla et al., 2024a).
@@ 2 BTSbot-nearby
We make some minor adjustments to criteria in these filters. The real/bogus score threshold using braai (Duev et al., 2019), ZTF’s deep learning model for filtering of non-astrophysical alerts, is increased from 0.3 (in the BTS alert fitler) to 0.7. This adjustment is made to more consistently reject bogus alerts which frequently occur atop very bright galactic nuclei. Such backgrounds which are common for nearby transients and alerts from these areas tend to be misclassified by braai. Based on Figure 9 of Duev et al. (2019), this should reduce the real/bogus false positive rate from 3.6% to 1.6%. We also make the rejection of candidates based on crossmatches to the Minor Planet Center (MPC) catalog more strict; a source with any alert within 2” of an MPC object will be rejected.
@@
The primary addition is filtering following host galaxy association. Alerts only pass the BTSbot-nearby filter if they are coincident with a NASA Extragalactic Database Local Volume Sample (NED-LVS; Cook et al., 2023, 2025) galaxy that has distance Mpc. An alert is considered coincident with a galaxy if its projected physical offset is 15 kpc, capped at a maximum angular offset of . The angular offset limit is imposed because the size of the selection region for very nearby galaxies ( Mpc) is otherwise extremely large and prone to selecting sources unrelated to the nearby galaxy. The 60 Mpc limit is motivated by the ZTF sensitivity: a 20.5 mag source at 60 Mpc corresponds to having an absolute magnitude of roughly mag, sufficient to capture SNe very early. An offset limit of 15 kpc selects the vast majority of CCSNe and SNe Ia (Fremling et al., 2020; Schulze et al., 2021) while limiting contamination from sources projected nearby but unrelated to a galaxy.
@@
Once a source is associated to a galaxy, the galaxy’s distance can be used to estimate the source’s absolute magnitude. BTSbot-nearby requires that the source have absolute magnitude in any ZTF filter of mag to reject variable stars and classical novae which are visible to ZTF in the most nearby galaxies (e.g., M31 and M33; Capaccioli et al., 1989). The source must also have at least one alert with BTSbot score greater than 0.5. This criterion is crucial for rejecting remaining contamination like non-SN stellar activity in very nearby galaxies, active galactic nuclei and cataclysmic variables projected over nearby galaxies, and more. The target must also not be coincident with a source on the Transient Name Server (TNS)333https://www.wis-tns.org that was reported more than 36 hours before the time of filtering. Next, we select only sources with non-detections within 3.5 days of the latest alert to ensure that those transients which are selected are very young. The final criteria considers the status of the source on ZTF’s first-party marshal Fritz444https://github.com/fritz-marshal/fritz (a SkyPortal instance; van der Walt et al., 2019; Coughlin et al., 2023); a trigger is prevented if there is (i) an existing spectroscopic follow-up request; (ii) a classification assigned; or (iii) a spectrum already present. The criteria involving the status of the source on Fritz are described in detail in Appendix B of Rehemtulla et al. (2024a). Once all criteria are met, follow-up requests can be triggered with configurable parameters.
@@ 2.1 Quality of BTSbot-nearby filtering relative to human scanning
We characterize the quality of the BTSbot-nearby filtering by assessing the completeness (or “recall”) and purity (or “precision”) of triggers and comparing the latency of automated follow-up with that of traditional, human-triggered follow-up. We pass ZTF public and partnership alerts from the main surveys between 1 October 2023 to 1 April 2025555The dates are determined by when the latest version of BTSbot was deployed into production and when this analysis was conducted. through the BTSbot-nearby filters and log which sources satisfy them and whether they are true positives (TPs) or FPs. False negatives (FNs) are identified by cross-matching with sources cataloged in BTS or CLU during the same time frame.
Statistics are calculated identically to how they were by Rehemtulla et al. (2024a). 77 sources satisfy the BTSbot-nearby filtering in this time frame; of these are genuine nearby transients, identified via cross-match with BTS/CLU catalogs and visual inspection. The FPs are strongly dominated by distant transients that are erroneously cross-matched to nearby galaxies but also include small numbers of active galactic nuclei (AGN) and cataclysmic variable stars (CVs). The count of FNs are definition dependent. In most cases, BTS/CLU transients were not selected by BTSbot-nearby filtering because their host galaxies do not appear in the NED catalog used. In some cases, the host galaxies had entries in NED but lacked either a spectroscopic redshift or a redshift-independent distance measurement. Transients which are very highly offset from their host represent a very small but astrophysically important population of FNs. These results suggest that that the host galaxy association mechanism used by BTSbot-nearby can still be improved.
@@
Figure 1 shows the distribution of spectroscopic follow-up latencies for nearby SNe, highlighting those involving BTSbot and that which was demonstrated by BTSbot-nearby for SN 2024jlf. Latencies for SNe with human-triggered follow-up clearly cluster around day. This matches expectations for this time-period because visual inspection and triggering of high-priority follow-up typically only occur the morning after observations. When this is the case and the follow-up instrument is co-located with the survey (e.g. ZTF and SEDM), follow-up tends to occur 24 hours after discovery. This amount of latency or more is reflected in the results of numerous infant SN studies including those which deal with UV follow-up; see, e.g., Figure 2 of Irani et al. (2024).
This is not always the case, however. Scanners in Europe are able to easily scan in near real-time because the Palomar night coincides with morning and afternoon in Europe. This has resulted in countless instances of early follow-up triggered by European scanners yielding valuable results (e.g., Yang et al., 2021; Sollerman et al., 2021). In fact, much of the follow-up effort supporting Bruch et al. (2021) and B23 was enabled by scanners in Europe. Because the sample used here is bounded by 1 January 2021, it does not contain any SNe in B23, but the latency statistics are representative for human scanning between January 2021 and October 2024.
@@ 3.1 Discovery and Classification
Before either TNS report, BTSbot identified ZTF24aaozxhx as a genuine bright ( mag) transient, saved it to an internal ZTF transient catalog, and triggered a high-priority follow-up request for photometry () and spectroscopy to SEDM at 07:53:40 (MJD 60458.33).666In the future, BTSbot-nearby discoveries will automatically be reported to TNS to facilitate community follow-up of these important targets. SEDM began observing ZTF24aaozxhx just 7 minutes later at 08:00:51; this observation occurred only 0.7 days after first light (see Sec. 3.3.1 for details). Logs of Fritz activity, where ZTF partners scan new transient candidates, suggest that no astronomer in ZTF had viewed ZTF24aaozxhx before the spectroscopic observation had concluded. Following this discovery, we collected multi-wavelength, high-cadence follow-up observations with numerous facilities to probe the evolution of SN 2024jlf.
@@ 3.3 Photometric observations
SN 2024jlf was observed in the UV (, , ) and near-UV () with the Ultraviolet/Optical Telescope (UVOT; Roming et al., 2005) onboard the Neil Gerhels Swift Obervatory (Swift; Gehrels et al., 2004). We processed all UVOT photometry of SN 2024jlf using a Python-based wrapper777https://github.com/charliekilpatrick/Swift_host_subtraction for photometry tools in heasoft v6.28 (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc). We optimally stacked each epoch and performed aperture photometry in each band using uvotsource.
@@ 3.4 Spectroscopic observations
We conducted a thorough spectroscopic follow-up campaign to monitor the evolution of SN 2024jlf out to 94.6 days after . Spectroscopic follow-up was conducted with numerous facilities: SEDM (Blagorodnova et al., 2018; Kim et al., 2022) on the P60 telescope (Cenko et al., 2006); the Spectrograph for the Rapid Acquisition of Transients (SPRAT; Piascik et al., 2014) on the robotic 2 m Liverpool Telescope (LT; Steele et al., 2004); the Alhambra Faint Object Spectrograph and Camera (ALFOSC)888https://www.not.iac.es/instruments/alfosc/ on the 2.56 m Nordic Optical Telescope (NOT); Binospec (Fabricant et al., 2019) on the 6.5 m MMT telescope; the Low Resolution Imaging Spectrometer (LRIS; Oke et al., 1995) on the W. M. Keck Observatory’s Keck I 10 m telescope; the Keck Cosmic Web Imager (KCWI; Martin et al., 2010; Morrissey et al., 2018) on the Keck II 10 m telescope; and the Wide-Field Spectrograph (WiFeS; Dopita et al., 2007, 2010) on the Australian National University 2.3 m (ANU) Advanced Technology Telescope (ATT) at Siding Spring Observatory. We also include the FLOYDS-N spectrum uploaded to TNS by Hosseinzadeh et al. (2024).
@@
The SEDM observations are reduced by the custom pySEDM package (Rigault et al., 2019). Spectra from other facilities are reduced with standard procedures (see Matheson et al. 2000). Reduction of MMT/Binospec and Keck/LRIS spectra uses pypeit (Prochaska et al., 2020); LT/SPRAT reduction uses the custom FRODOSpec pipeline (Barnsley et al., 2012); NOT/ALFOSC reduction uses a custom pypeit-based script;999https://gitlab.com/steveschulze/pypeit_alfosc_env Keck/KCWI reduction uses the official data reduction pipeline;101010https://kcwi-drp.readthedocs.io/ and ANU/WiFeS reduction uses PyWiFeS (see Childress et al., 2014; Carr et al., 2024).111111https://github.com/PyWiFeS/pipeline Our spectroscopic observations are summarized in Table 3.
@@
Figure 6 shows our full spectral series of SN 2024jlf. We find that SN 2024jlf exhibits features and evolution typical of SNe IIP: a mostly featureless blue continuum followed by prominent P-Cygni lines of the hydrogen Balmer series and later of He. A broad H profile first emerges in the day SPRAT spectrum. H and bluer Balmer series lines become visible shortly afterwards; the Binospec spectrum at days shows P-Cygni profiles in H, H, H, H, and possibly H. He I is also visible in a P-Cygni profile starting in the day spectrum. Fe II absorption and the Ca near-infrared triplet are visible starting around three weeks after first light. Model spectra, shown as gray lines, are discussed in Section 4.1.
@@
Figure 7 shows the regions around main spectral lines in the early spectra of SN 2024jlf which exhibit features characteristic of SNe II undergoing CSM interaction: He II 4686 and the ledge, C IV 5801, 5812, and H. We also possibly detect narrow H emission in the d FLOYDS spectrum. He II is very common in flashing SNe and its presence is often used to define a SN as flashing (e.g., Khazov et al. 2016; Bruch et al. 2021; B23). The ledge-shaped feature around 4450–4700 Å, visible in the first six epochs of spectroscopy, is typically interpreted as a sign of CSM interaction and has been seen in many interacting SNe including SN 2013fs (Yaron et al., 2017) and others in the samples of B23 and JG24. Its origin is not precisely known although it has been attributed to (i) a blend of high-ionization lines including those of carbon, nitrogen, and oxygen (e.g., Soumagnac et al., 2020; Bruch et al., 2021; Pearson et al., 2023); (ii) high-velocity H (e.g., Pastorello et al., 2006); and (iii) blueshifted He II 4686 (e.g., Dessart et al., 2017; Bullivant et al., 2018; Chugai & Utrobin, 2023). See Pearson et al. (2023) for further discussion and references on the ledge feature.